# Periodic classical trajectories and quantum scars in many-spin systems

### v2

@@ VI.2 Quantum scars in finite-size spin clusters
We searched for the presence of quantum scars by numerically computing the eigenstates of the spin chains, and then computing for each eigenstate the half-chain entanglement entropy $\mathcal{E}$ between the sites from 1 to $L/2$ (rounded down for odd $L$) and the rest of the chain. The entanglement entropy is a popular choice for monitoring eigenstate thermalisation; other physical observables could have been used as well. ETH implies that, in the dense part of the spectrum, $\mathcal{E}$ as a function of eigenstate’s energy $E_{n}$ changes only very little from one state to the next. A large departure of $\mathcal{E}$ from the value typical of the adjacent-in-energy eigenstates is then an indication of a quantum scar.
Figure 18: Quantum-scar diagnostics for quantum chains of lengths $L$ (indicated above each column of plots) for spins S equal to 1/2, 1, 3/2 and 2 (indicated on the left of the plots). The interaction constant in all cases is $\tilde{J}=1.76$. For each value of $S$, the upper line of plots represents normalized half-chain entanglement entropy $\mathcal{E}$ for translationally invariant (zero momentum) Hamiltonian eigenstates $|E_{n}\rangle$ as a function of eigenenergies $E_{n}$. The outliers of $\mathcal{E}$ are identified with quantum scars. The bottom line of plots for each $S$ represents the overlaps $|\langle\Psi^{\text{up}}|E_{n}\rangle|^{2}$ of fully polarized “spins-up” state $|\Psi^{\text{up}}\rangle$ with $|E_{n}\rangle$. Magenta points in the plots $\mathcal{E}(E_{n})$ correspond to the top five eigenstates having the largest overlap values $|\langle\Psi^{\text{up}}|E_{n}\rangle|^{2}$ indicated by magenta bars in the respective overlap plots. The vertical axis of each panel is scaled such the maximum projection of the data points on that axis is equal to 1.
The computed entanglement entropy is plotted in Fig. 18 for the chains of spins 1/2, 1, 3/2 and 2. These plots reveal the picture entirely consistent with the expectations based on observing or not observing the slowdown of thermalisation for the initial state $|\Psi^{\text{up}}\rangle$ — see Fig. 16. Namely, quantum-scar eigenstates are prominently present for the chains of spins 3/2 and 2 as indicated by the outliers of $\mathcal{E}(E_{n})$ around the center of the spectrum, while the chains of spins 1/2 exhibit no signs of quantum scars, and spin-1 chains show rather weak indications thereof.
@@
Figure 19: Spherical map of the overlap $|\langle E_{n}|\Psi(\theta,\phi)\rangle|^{2}$ between an energy eigenstate $|E_{n}\rangle$ and the state $|\Psi(\theta,\phi)\rangle$ of all spins maximally polarised along the direction $\mathbf{n}=(\sin{\theta}\cos{\phi},\sin{\theta}\sin{\phi},\cos{\theta})$ for a chain of spins 2 with ${\tilde{J}=1.76}$ and $L=6$. The red line represents the periodic trajectory in the classical limit. (a) Overlap for the most prominent quantum scar eigenstate with $n=927$ (see the relevant panel in Fig. 18). (b) Overlap for a nearby generic energy eigenstate with $n=925$. ($n=926$ corresponds to a weaker quantum scar.)
In order to visualise the above connection, we present in Fig. 19(a) a spherical map of the overlap $|\langle E_{n}|\Psi(\theta,\phi)\rangle|^{2}$ between the most prominent quantum-scar eigenstate $|E_{n}\rangle$ for a chain of spins 2 and the state
@@
where every spin is fully polarised along the direction $\mathbf{n}=(\sin{\theta}\cos{\phi},\sin{\theta}\sin{\phi},\cos{\theta})$ determined by the standard spherical angles $\theta$ and $\phi$. The concentration of this projection around the corresponding classical trajectory is evident there. For comparison, Fig. 19(b) shows the overlap with $|\Psi(\theta,\phi)\rangle$ for an adjacent generic energy eigenstate: In the latter case, the overlap exhibits much smaller values at the level of statistical fluctuations in a high-dimensional Hilbert space.
The tendency of larger quantum spins $S$ to exhibit more pronounced evidence of quantum scarring can be understood as follows. For a billiard-like quantum many-body scar to be visible, the quantum dynamics should be close enough to a semiclassical limit, in which case, the difference between a quantum scar and a generic eigenstate is as dramatic as the difference between short periodic and ergodic classical trajectories. Without proximity to the semiclassical limit, fluctuations associated with strong quantum interference effects are, presumably, making it difficult to distinguish generic from nongeneric eigenstates. In our case, Fig. 16 illustrates that the accuracy of the semiclassical approximation for quantum spin dynamics indeed improves quickly with the growth of $S$.
The presented evidence of quantum scars is limited only to spin chains of finite size. It remains an intriguing question, whether the quantum-scar eigenstates remain present in the thermodynamic limit $L\to\infty$ for the chains of spins $S\geq 3/2$. On the one hand, periodic classical trajectories should become more unstable as $L$ grows, because their overall instability is characterised by the sum of all positive Lyapunov exponents (Kolmogorov-Sinai entropy), which should grow proportionally to $L$ [120]; hence the lifetime of coherent semiclassical quantum wave packets near periodic trajectories should decrease, thereby suppressing quantum interference behind the formation of quantum scars. On the other hand, the increasing instability of many-spin classical periodic trajectories does not necessarily exclude the existence of exponentially few quantum scars out of $(2S+1)^{L}$ energy eigenstates. One can further argue that the atypical slowdown of thermalisation, which, as shown in the preceding subsection, remains present in the thermodynamic limit, indicates that, independent of the size of the system, there is an atypical sector of initial conditions in the many-spin Hilbert space, and, while a randomly selected wave function from that sector is not an eigenstate, there is a chance that a superposition of wave functions exclusively belonging to the atypical sector does constitute an eigenstate, and this would be a quantum scar remaining present in the thermodynamic limit.
@@ VI.3 Quantum separatrix
The existence of the above two quantum regimes led us to further explore, whether the notion of a separatrix can be defined in the quantum case. Indeed, beyond the visual impression of the relaxation crossover, the quantum separatrix value for $\tilde{J}$ can be accurately identified as follows. We note that, in the classical limit, the separatrix appears at a saddle point of the energy function $E(S_{x},S_{y},S_{z})$ describing the periodic trajectories, where each spin has the same orientation ($S_{x},S_{y},S_{z}$). The classical phase-space volume available for the motion around that saddle point is supposed to have a Van Hove singularity as a function of energy. In the quantum case, it is thus reasonable to expect that the number of the relevant quantum energy eigenstates $|E_{n}\rangle$ has a maximum around the quantum separatrix for finite systems and, possibly, a singularity in the thermodynamic limit. As a measure of that number, we use the participation ratio $\mathcal{P}$ of eigenstates $\left\{|E_{n}\rangle\right\}$ in the initial state $|\Psi^{\text{up}}\rangle$:
@@ VII Summary and discussion
(C1) The largest Lyapunov exponents of the periodic trajectories exhibit nontrivial nonmonotonic dependence on the length of spin chains presented in Fig. 4. We were able to find the functional form of this dependence, Eq.(12), on the basis of the insight that the Lyapunov vectors for the Lyapunov instabilities around translationally invariant periodic trajectories constitute irreducible representations of the translational symmetry group with well-defined wave vectors.
@@
(C2) Our particularly surprising finding is that the periodic trajectories of rather long spin chains – up to 59 spins in Fig. 4(c) – can be Lyapunov stable. Both our numerical experience and the proposed analytical form (12) of $\lambda_{p}(L)$ suggest that the maximum length of chains with stable periodic trajectories is limited for a given value of the coupling constant $J$. However, as the value of $J$ increases, so does the maximum length of the stable periodic motion, i.e. one may, possibly, be able to make the periodic trajectories stable for a spin chain of any given length by choosing large enough $J$. Another counter-intuitive aspect of the same observation is that the periodic motion in a system of a large size can be perfectly stable, while it is unstable in a much smaller system with the same coupling constants.
@@
(C5) We have shown in Section V that, once a Lyapunov instability develops around a periodic trajectory, the system generally enters a transient nearly quasiperiodic regime. That regime is typically characterised by an interplay of two spatial Fourier harmonics: one corresponding to the original periodic motion and the other to the leading Lyapunov instability. The nearly quasiperiodic character of the motion is well defined, because the two Fourier harmonics are described by a four-dimensional phase space, where KAM tori of quasi-periodic motion are known to constrain stochastic trajectories. We have identified three mechanisms that destroy this regime. The first two of them are relatively fast — they are associated with additional Lyapunov instabilities involving other spatial Fourier harmonics. The third mechanism is extremely slow: it is the Arnold diffusion that sets in those rather exceptional cases, where no additional Lyapunov instabilities are available. This observation is rather remarkable, given that Arnold diffusion is generaly expected for nearly integrable Hamiltonians, while in the present case the Hamiltonian is strongly nonintegrable. Yet Arnold diffusion becomes relevant due to two factors: (i) weak nonintegrability of the dynamics in the close vicinity of the considered periodic trajectories and (ii) the translational invariance of these trajectories that affords the possibility that all spatial Fourier harmonics beside the above mentioned two remain stable.
@@
(Q1.A) The first such signature is the anomalous slowdown of the thermalisation process starting from the initial quantum states corresponding to the classical periodic trajectories. This signature is shown to remain present in the thermodynamic limit.
(Q1.B) The second signature is that energy eigenfunctions dominating the expansion of the initial quantum states associated with the classical periodic motion exhibit nonconformance with the eigenstate thermalisation hypothesis. While the initial proposal of many-body quantum scars[39, 105] focused on anomalous energy eigenstates in spin-1/2 systems without a clear classical counterpart, here we demonstrated the viability of the original definition of quantum scars[114] as the eigenstates corresponding to periodic trajectories in the classical limit.
In this work, we numerically identified the quantum-scar states only in the finite-size systems. The survival of these states in the thermodynamic limit remains an interesting mathematical question. Let us, however, remark that the observed anomalous slowdown of thermalisation in the thermodynamic limit would constitute, perhaps, the most important practical consequence of the existence of the quantum scars. At the same time, our result that finite systems possess quantum-scar eigenstates which, on the one hand, are linked to periodic classical trajectories, and, on the other hand, control the thermalisation slowdown, clearly indicates that the mechanism behind the slowdown is the quantum interference along the short periodic classical trajectories. Since the slowdown remains present in the thermodynamic limit, the interference effect along the short periodic trajectories should also survive the thermodynamic limit. One can, therefore, argue that the thermalisation slowdown together with the presence of quantum scar eigenstates for finite systems can be taken as a primary definition of quantum scars in many-body systems irrespective of whether eigenstates violating ETH actually exist in the thermodynamic limit.
As mentioned in Section II.2, our findings are broadly consistent with those of Ref. [113] and together build the case that quantum many-body systems generically exhibit quantum scars associated with short periodic trajectories realizing irreducible representations of the translational symmetry group with a well-defined wave vector.
@@
In the present work, we have not looked for the quantum signatures of the detected Arnold diffusion, but it remains an interesting direction to explore both in terms of quantum localization properties on the Arnold web[49] and in terms of possible hybridization of eigenstates[50] between ergodic and non-ergodic regions of the phase space.
We, finally, mention, that, according to the proposal of Ref. [156], the connection between quantum scars and unstable periodic classical trajectories can be exploited to exercise dynamical control over large clusters of quantum particles.
@@ Acknowledgements.
Appendix A Numerical calculation of Lyapunov exponents

### v3

@@ VI.2 Quantum scars in finite-size spin clusters
We searched for the presence of quantum scars by numerically computing the eigenstates of the spin chains, and then computing for each eigenstate the half-chain entanglement entropy $\mathcal{E}$ between the sites from 1 to $L/2$ (rounded down for odd $L$) and the rest of the chain. The entanglement entropy is a popular choice for monitoring eigenstate thermalization; other physical observables could have been used as well. ETH implies that, in the dense part of the spectrum, $\mathcal{E}$ as a function of eigenstate’s energy $E_{n}$ changes only very little from one state to the next. A large departure of $\mathcal{E}$ from the value typical of the adjacent-in-energy eigenstates is then an indication of a quantum scar.
Figure 18: Quantum-scar diagnostics for quantum chains of lengths $L$ (indicated above each column of plots) for spins S equal to 1/2, 1, 3/2 and 2 (indicated on the left of the plots). The interaction constant in all cases is $\tilde{J}=1.76$. For each value of $S$, the upper line of plots represents normalized half-chain entanglement entropy $\mathcal{E}$ for translationally invariant (zero momentum) Hamiltonian eigenstates $|E_{n}\rangle$ as a function of eigenenergies $E_{n}$. The outliers of $\mathcal{E}$ are identified with quantum scars. The bottom line of plots for each $S$ represents the overlaps $|\langle\Psi^{\text{up}}|E_{n}\rangle|^{2}$ of fully polarized “spins-up” state $|\Psi^{\text{up}}\rangle$ with $|E_{n}\rangle$. Magenta points in the plots $\mathcal{E}(E_{n})$ correspond to the top five eigenstates having the largest overlap values $|\langle\Psi^{\text{up}}|E_{n}\rangle|^{2}$ indicated by magenta bars in the respective overlap plots. The vertical axis of each panel is scaled such the maximum projection of the data points on that axis is equal to 1. The trends exhibited in this figure are discussed in Section VI.2.
The computed entanglement entropy is plotted in Fig. 18 for the chains of spins 1/2, 1, 3/2 and 2. These plots reveal the picture entirely consistent with the expectations based on observing or not observing the slowdown of thermalization for the initial state $|\Psi^{\text{up}}\rangle$ — see Fig. 16. Namely, quantum-scar eigenstates are prominently present for the chains of spins 3/2 and 2 as indicated by the outliers of $\mathcal{E}(E_{n})$ around the center of the spectrum, while the chains of spins 1/2 exhibit no signs of quantum scars, and spin-1 chains show rather weak indications thereof.
@@
Figure 19: [Color online] Spherical map of the overlap $|\langle E_{n}|\Psi(\theta,\phi)\rangle|^{2}$ between an energy eigenstate $|E_{n}\rangle$ and the state $|\Psi(\theta,\phi)\rangle$ of all spins maximally polarised along the direction $\mathbf{n}=(\sin{\theta}\cos{\phi},\sin{\theta}\sin{\phi},\cos{\theta})$ for a chain of spins 2 with ${\tilde{J}=1.76}$ and $L=6$. The red line represents the periodic trajectory in the classical limit. (a) Overlap for the most prominent quantum scar eigenstate with $n=927$ (see the relevant panel in Fig. 18). (b) Overlap for a nearby generic energy eigenstate with $n=925$. ($n=926$ corresponds to a weaker quantum scar.)
In order to visualize the above connection, we present in Fig. 19(a) a spherical map of the overlap $|\langle E_{n}|\Psi(\theta,\phi)\rangle|^{2}$ between the most prominent quantum-scar eigenstate $|E_{n}\rangle$ for a chain of spins 2 and the state
@@
where every spin is fully polarized along the direction $\mathbf{n}=(\sin{\theta}\cos{\phi},\sin{\theta}\sin{\phi},\cos{\theta})$ determined by the standard spherical angles $\theta$ and $\phi$. The concentration of this projection around the corresponding classical trajectory is evident there. For comparison, Fig. 19(b) shows the overlap with $|\Psi(\theta,\phi)\rangle$ for an adjacent generic energy eigenstate: In the latter case, the overlap exhibits much smaller values at the level of statistical fluctuations in a high-dimensional Hilbert space.
The tendency of larger quantum spins $S$ to exhibit more pronounced evidence of quantum scarring can be understood as follows. In order to have a quantum scar distinctively localized around a periodic classical trajectory (as in the case of billiards[115]), the quantum dynamics should be close enough to a semiclassical limit in the first place, such that a quantum wave packet initiated around that trajectory does not broaden too much before the trajectory closes upon itself. When this condition is fulfilled, the difference between a quantum scar and a generic eigenstate is as dramatic as the difference between short periodic and ergodic classical trajectories. Without proximity to the semiclassical limit, fluctuations associated with strong quantum interference effects are, presumably, making it difficult to distinguish generic from nongeneric eigenstates. In our case, Fig. 16 illustrates that the accuracy of the semiclassical approximation for quantum spin dynamics indeed improves quickly with the growth of $S$.
The presented evidence of quantum scars is limited only to spin chains of finite size. It remains an intriguing question, whether the quantum-scar eigenstates remain present in the thermodynamic limit $L\to\infty$ for the chains of spins $S\geq 3/2$. On the one hand, periodic classical trajectories should become more unstable as $L$ grows, because their overall instability is characterised by the sum of all positive Lyapunov exponents (Kolmogorov-Sinai entropy), which should grow proportionally to $L$ [126]; hence the lifetime of semiclassical quantum wave packets made of coherent spin states[165] near periodic trajectories should decrease, thereby suppressing quantum interference behind the formation of quantum scars. On the other hand, the increasing instability of many-spin classical periodic trajectories does not necessarily exclude the existence of a few quantum scars out of exponentially many, i.e. $(2S+1)^{L}$, energy eigenstates. One can further note that the atypical slowdown of thermalization seen in Figs. 16(b-d) and 17 is very likely to remain present in the thermodynamic limit, and this would indicate that, independent of the size of the system, there is an atypical sector of initial conditions in the many-spin Hilbert space. Although a randomly selected wave function from that sector is not an eigenstate, we conjecture that at least a few quantum-scar eigenstates consisting of the superposition of wave functions predominantly belonging to the above atypical sector do remain present in the thermodynamic limit.
@@ VI.3 Quantum separatrix
The existence of the above two quantum regimes led us to further explore, whether the notion of a separatrix can be defined in the quantum case. Indeed, beyond the visual impression of the relaxation crossover, the quantum separatrix value for $\tilde{J}$ can be accurately identified as follows. We note that, in the classical limit, the separatrix appears at a saddle point of the Hamiltonian function $\mathcal{H}_{\text{p}}(S^{x}_{\text{p}},S^{y}_{\text{p}},S^{z}_{\text{p}})$ given by Eq.(9) and describing the periodic trajectories, where each spin has the same orientation $(S^{x}_{\text{p}},S^{y}_{\text{p}},S^{z}_{\text{p}})$. The classical phase-space volume available for the motion around that saddle point has a Van Hove singularity as a function of energy. In the quantum case, it is thus reasonable to expect that the number of the relevant quantum energy eigenstates $|E_{n}\rangle$ has a maximum around the quantum separatrix for finite systems and, possibly, a singularity in the thermodynamic limit. As a measure of that number, we use the participation ratio $\mathcal{P}$ of eigenstates $\left\{|E_{n}\rangle\right\}$ in the initial state $|\Psi^{\text{up}}\rangle$:
@@ VII Summary and discussion
(C1) The largest Lyapunov exponents of the periodic trajectories exhibit nontrivial nonmonotonic dependence on the length of spin chains presented in Fig. 4. We were able to find accurate quantitative approximation (13) for the functional form of this dependence on the basis of the insight that the Lyapunov vectors for the Lyapunov instabilities around translationally invariant periodic trajectories constitute irreducible representations of the translational symmetry group with well-defined wave vectors.
@@
(C2) Our particularly surprising finding is that the periodic trajectories of rather long spin chains – up to 59 spins in Fig. 4(c) – can be Lyapunov stable. (“Rather long” in the sense of the phase space dimension $2L$ being much larger than 1.) Both our numerical experience and the proposed analytical form (13) of $\lambda_{p}(L)$ suggest that the maximum length of chains with stable periodic trajectories is limited for a given value of the coupling constant $J$. However, as the value of $J$ increases, so does the maximum length of the stable periodic motion, i.e. one may, possibly, be able to make the periodic trajectories stable for a spin chain of any given length by choosing large enough $J$. Another counter-intuitive aspect of the same observation is that the periodic motion in a system of a large size can be perfectly stable, while it is unstable in a much smaller system with the same coupling constants.
@@
(C5) We have shown in Section V that, once a Lyapunov instability develops around a periodic trajectory, the system generally enters a transient nearly quasiperiodic regime (meaning a stochastic regime close to a quasiperiodic motion). That regime is typically characterised by an interplay of two spatial Fourier harmonics: one corresponding to the original periodic motion and the other to the leading Lyapunov instability. The nearly quasiperiodic character of the motion arises, because the two Fourier harmonics are described by a four-dimensional phase space, where KAM tori of quasiperiodic motion divide energy shells into impenetrable parts, and, hence, two nearby KAM tori can “squeeze” a stochastic trajectory and thus force the nearly quasiperiodic regime. We have identified three mechanisms that destroy this regime. The first two of them are relatively fast (in comparison with the third one) — they are associated with additional Lyapunov instabilities involving other spatial Fourier harmonics. The third mechanism is extremely slow: it is the Arnold diffusion that sets in those rather exceptional cases, where no additional Lyapunov instabilities are available. This observation is rather remarkable, given that Arnold diffusion is generaly expected for nearly integrable Hamiltonians, while in the present case the Hamiltonian is strongly nonintegrable. Yet Arnold diffusion becomes relevant due to two factors: (i) weak nonintegrability of the dynamics in the close vicinity of the considered periodic trajectories and (ii) the translational invariance of these trajectories that affords the possibility that all spatial Fourier harmonics beside the above mentioned two remain stable.
@@
(Q1.A) The first such signature is the anomalous slowdown of the thermalization process starting from the initial quantum states corresponding to the classical periodic trajectories. This signature is shown to remain present in the thermodynamic limit.
(Q1.B) The second signature is that energy eigenfunctions dominating the expansion of the initial quantum states associated with the classical periodic motion exhibit nonconformance with the eigenstate thermalization hypothesis. While the initial proposal of many-body quantum scars[39, 106] focused on anomalous energy eigenstates in spin-1/2 systems without a clear classical counterpart, here we demonstrated the viability of the original definition of quantum scars[115] as the eigenstates corresponding to periodic trajectories in the classical limit.
In this work, we numerically identified the quantum-scar states only in the finite-size systems. The survival of these states in the thermodynamic limit remains an interesting mathematical question. Let us, however, remark that the observed anomalous slowdown of thermalization in the thermodynamic limit would constitute, perhaps, the most important practical consequence of the existence of the quantum scars. At the same time, our result that finite systems possess quantum-scar eigenstates which, on the one hand, are linked to periodic classical trajectories, and, on the other hand, control the thermalization slowdown, indicates that the mechanism behind the formation of quantum scars, namely, the quantum interference along short periodic classical trajectories[115], is also the mechanism behind the thermalization slowdown. Since the slowdown remains present in the thermodynamic limit, the interference effect along the short periodic trajectories should also survive the thermodynamic limit. One can, therefore, argue that the thermalization slowdown together with the presence of quantum scar eigenstates for finite systems can be taken as a primary definition of quantum scars in many-body systems irrespective of whether eigenstates violating ETH actually exist in the thermodynamic limit.
As mentioned in Section II.2, our findings are broadly consistent with those of Refs. [120, 121, 122, 114] and together build the case that translationally invariant quantum many-body systems generically exhibit quantum scars associated with short periodic trajectories realizing irreducible representations of the translational symmetry group with a well-defined wave vector.
@@
Let us now make two final remarks connecting this work to the recent literature: (i) We have not looked for the quantum signatures of the detected Arnold diffusion, but it remains an interesting direction to explore both in terms of quantum localization properties on the Arnold web[49] and in terms of possible hybridization of eigenstates[50] between ergodic and non-ergodic regions of the phase space. (ii) According to the proposal of Ref. [170], the connection between quantum scars and unstable periodic classical trajectories can be exploited to exercise dynamical control over large clusters of quantum particles.
@@ Acknowledgements.
Appendix A Periodicity of one-spin classical trajectories $\mathbf{S}_{\text{p}}(t)$ and the transition from librations to rotations
In order to follow the transition from librations to rotations analytically and simultaneously demonstrate the generic periodicity of the one-spin orbits $\mathbf{S}_{\text{p}}(t)$, we note that energy $E=0$ corresponding to our initial conditions “all-spins-up”, implies a constraint on the time evolution of $\mathbf{S}_{\text{p}}(t)$, which, given the Hamiltonian (9), can be rewritten as
$\left(S^{x}_{\text{p}}-\frac{1}{2}\frac{h}{J}\right)^{2}+2\left(S^{y}_{\text{p}}-\frac{1}{4}\frac{h}{J}\right)^{2}=\frac{3}{8}\left(\frac{h}{J}\right)^{2}.$ (31)
This equation defines an ellipse in the $(S^{x}_{\text{p}},S^{y}_{\text{p}})$-plane. At the same time, the spin normalization condition implies:
$(S^{x}_{\text{p}})^{2}+(S^{y}_{\text{p}})^{2}=1-(S^{z}_{\text{p}})^{2}\leq 1$ (32)
i.e., the projection of the spin on the $(S^{x}_{\text{p}},S^{y}_{\text{p}})$ plane is bounded by a unit circle. As illustrated in Fig. 22, the intersection of the ellipse (31) and the unit circle (32) determines the type of motion of the spin on the unit sphere as follows:
(i) A libration corresponds to the case where the ellipse (31) intersects the unit circle at two points associated with the spin transitioning back and forth between the hemispheres $S^{z}_{\text{p}}>0$ and $S^{z}_{\text{p}}<0$. The elliptic segment truncated by the unit circle is a projection by the libration trajectory onto the $(S^{x}_{\text{p}},S^{y}_{\text{p}})$-plane. Each point on this projection is associated with two points on the libration trajectory and thus is being passed twice during the libration period.
(ii) A rotation is the case where the ellipse lies entirely within the unit circle without intersecting it. In this case, the motion is limited to either hemispheres $S^{z}_{\text{p}}>0$ or $S^{z}_{\text{p}}<0$. Correspondingly, there are two disconnected rotation trajectories having the ellipse as their $(S^{x}_{\text{p}},S^{y}_{\text{p}})$-projection.
(iii) A separatrix emerges, when the ellipse touches the unit circle at one point, which then becomes the fixed point.
In order to establish the correspondence between the above regimes and the value of $J/h$, we note that the ellipse (31) must pass through the point $(S^{x}_{\text{p}},S^{y}_{\text{p}})=(0,0)$ for any $J/h$. Small values of $J/h$ imply large linear size of the ellipse, thereby guaranteeing two crossing points with the unit circle and hence the libration regime, while large values of $J/h$ imply that the ellipse is too small and thus cannot cross the unit circle, which means the rotation regime.
We, finally, note that, for a broader class of spin Hamiltonians, one can imagine a larger variety of situations, such as, e.g., the ellipse crossing the circle at four points and still implying a pair of rotation trajectories, or the ellipse touching the unit circle at two points and implying heteroclinic non-periodic trajectories. In any case, the explicit formula (31) for the $(S^{x}_{\text{p}},S^{y}_{\text{p}})$-projection together with $S^{z}_{\text{p}}=\pm\sqrt{1-(S^{x}_{\text{p}})^{2}-(S^{y}_{\text{p}})^{2}}$ implies that periodic trajectories $\mathbf{S}_{\text{p}}(t)$ represent a generic case, while the non-periodic ones – homoclinic or heteroclinic - are also possible (as anticipated by the Poincaré-Bendixson theorem[127, 128]) but only as an exception.
Figure 22: Projections of one-spin trajectories $\mathbf{S}_{\text{p}}(t)$ onto the $(S^{x}_{\text{p}},S^{y}_{\text{p}})$-plane for $h=1$ and different values of $J$ in the Hamiltonian (9). Colored lines are the ellipses (or a part thereof) defined by Eq. (31) - among them: blue solid elliptic segment is a projection of libration ($J=0.76$), green solid ellipse is a projection of rotation ($J=1.76$), and brown dashed ellipse is a projection of the separatrix ($J=J^{*}\approx 1.15$). Black line is the unit circle defined by Eq. (32). The three plotted projections correspond to the three examples shown in Fig. 2.
Appendix B Upper bound on the largest Lyapunov exponent for classical spin dynamics
Here we derive the upper bound on the largest Lyapunov exponent for the classical spin chain considered in this work. The derivation is trivially generalizable to any spin lattice with finite number of interacting neighbors per spin: it amounts to an explicit implementation of the steps of the general argument given in Ref.[126] that the maximum Lyapunov exponent for classical spin dynamics must be finite.
The largest Lyapunov exponent defined by Eq.(10) for any trajectory, periodic or ergodic, can be computed as the time-average of the instantaneous “local stretching rates” along the trajectory:
[…]
where
$\lambda_{\text{inst}}(t)\equiv\frac{|\delta\dot{\mathcal{S}}(t)|}{|\delta\mathcal{S}(t)|}$ (34)
and $T$ is a sufficiently long integration time for an ergodic trajectory or a period for a periodic trajectory. Therefore,
$\lambda_{\text{max}}\leq\text{max}[\lambda_{\text{inst}}].$ (35)
Below we obtain a constraint from above on $\text{max}[\lambda_{\text{inst}}]$.
The linearization of the spin dynamics based on Eqs.(6,7) relatively to the small deviations of spin projections
$\delta\mathcal{S}(t)\equiv[\delta\mathbf{S}_{1}(t),\dots,\delta\mathbf{S}_{L}(t)]$ (36)
gives
$\delta\dot{\mathbf{S}}_{i}=\delta\mathbf{H}_{i}\times\mathbf{S}_{i}+\mathbf{H}_{i}\times\delta\mathbf{S}_{i},$ (37)
where $\mathbf{H}_{i}$ is given by Eq.(7), and
${\delta\mathbf{H}}_{i}=\left(\begin{array}[]{c}-J\ \delta S_{i-1}^{x}-J\ \delta S_{i+1}^{x}\\[4.0pt] -2J\ \delta S_{i-1}^{y}-2J\ \delta S_{i+1}^{y}\\[4.0pt] 0\end{array}\right)$ (38)
After Eq.(38) is substituted into Eq.(37), the result can be represented as
$\delta\dot{S}_{i}^{\alpha}=\sum_{j}^{\text{NN}(i)}A_{ij}^{\alpha\beta}\delta S_{i}^{\beta},$ (39)
where $\text{NN}(i)$ means that the sum is taken over the nearest neighbors of spin $\mathbf{S}_{i}$, and $A_{ij}^{\alpha\beta}$ is the tensor determined by the coupling constants and the instantaneous projections of $\mathbf{S}_{i}$ and its two neighbors. Since each projection of each spin has the maximum value 1, the maximum value of an element in this tensor can be constrained for our system by examining the coefficients in the explicit expressions for the individual projections of $\delta\dot{\mathbf{S}}_{i}$, which, in the case of Eq.(37), gives
$A_{\text{max}}\equiv\text{max}[|A_{ij}^{\alpha\beta}|]\leq 4|J|+|h|.$ (40)
After we collect all spin deviations into one $3L$-dimensional vector $\delta\mathcal{S}\equiv[\delta S^{x}_{1},\delta S^{y}_{1},\delta S^{z}_{1},\dots,\delta S^{x}_{L},\delta S^{y}_{L},\delta S^{z}_{L}]$, the counterpart of Eq.(39) reads
$\delta\dot{\mathcal{S}}=\mathcal{A}\cdot\delta\mathcal{S},$ (41)
where $\mathcal{A}$ is an $3L\times 3L$ tensor collecting all the entries of tensors $A_{ij}^{\alpha\beta}$. Below we also use variable $D=3L$ for the dimension of the extended phase space, and calligraphic variables $\delta\mathcal{S}_{n}$ and $\mathcal{A}_{mn}$ for the components of $\delta\mathcal{S}$ and $\mathcal{A}$. We note that, for any spin chain with nearest neighbor spin-spin interaction, each line or column of the tensor $\mathcal{A}$ cannot contain more than $K=9$ non-zero entries $\delta\mathcal{A}_{mn}$ originating from the three projections of each spin plus six projections of the two neighbors coupled to that spin.
The expression (34) for the local stretching rate can now be rewritten as
$\lambda_{\text{inst}}=\frac{|\mathcal{A}\delta\mathcal{S}|}{|\delta\mathcal{S}|},$ (42)
which implies that
$\lambda_{\text{inst}}^{2}=\frac{|\mathcal{A}\,\delta\mathcal{S}|^{2}}{|\delta\mathcal{S}|^{2}}=\frac{(\mathcal{A}_{11}\delta\mathcal{S}_{1}+\dots+\mathcal{A}_{1D}\delta\mathcal{S}_{D})^{2}+\dots+(\mathcal{A}_{D1}\delta\mathcal{S}_{1}+\dots+\mathcal{A}_{DD}\delta\mathcal{S}_{D})^{2}}{\delta\mathcal{S}_{1}^{2}+\dots+\delta\mathcal{S}_{D}^{2}}.$ (43)
We now proceed with the chain of inequalities for each term in the numerator of the right-hand-side (RHS) of Eq.(43)
$(\mathcal{A}_{n1}\delta\mathcal{S}_{1}+\dots+\mathcal{A}_{nD}\delta\mathcal{S}_{D})^{2}\leq\left(|\mathcal{A}_{n1}|\ |\delta\mathcal{S}_{1}|+\dots+|\mathcal{A}_{nD}|\ |\delta\mathcal{S}_{D}|\right)^{2}\leq A_{\text{max}}^{2}\left(\sum_{m(n)}^{K}|\delta\mathcal{S}_{m}|\right)^{2}\leq A_{\text{max}}^{2}K\sum_{m(n)}^{K}|\delta\mathcal{S}_{m}|^{2},$ (44)
where the notation $m(n)$ implies that the sum is taken only over the $K$ projections $|\delta\mathcal{S}_{m}|$ that were multiplied by those tensor elements $|\mathcal{A}_{nm}|$ that are allowed to have non-zero values. (Once our notations are traced back, a fixed value of index $n$ implies a certain projection $x$, $y$, or $z$ for one of our $L$ spins; the above sum then extends over $|\delta\mathcal{S}_{m}|$ corresponding to the three projections of that spin plus the six projections of its two coupled neighbors.) Once the inequality (44) is substituted for each term in the numerator of the RHS of Eq.(43), we further note that each squared projection $|\delta\mathcal{S}_{m}|^{2}$ of the vector $\delta\mathcal{S}$ can appear in the numerator maximum $K$ times, which, therefore, gives
$\lambda_{\text{inst}}^{2}\leq\frac{A_{\max}^{2}\ K^{2}\ (\delta S_{1}^{2}+\dots+\delta S_{M}^{2})}{\delta S_{1}^{2}+\dots+\delta S_{M}^{2}}=A_{\max}^{2}K^{2}.$ (45)
The inequalities (35) and (45) imply the upper bound
$\lambda_{\text{max}}\leq A_{\max}K$ (46)
for the largest Lyapunov exponent of both ergodic and periodic phase-space trajectories.
While the final inequality (46) firmly establishes the existence of the upper bound on $\lambda_{\text{max}}$, it is likely that the chain of inequalities leading to it can be made more restrictive, which would reduce that bound. We cannot provide any example where the value of $\lambda_{\text{max}}$ would actually reach $A_{\max}K$.
Appendix C Numerical calculation of Lyapunov exponents